A tennis ball is dropped from 1.18 m above the ground. It rebounds to a height of 1.05 m. With what velocity does it hit the ground? The acceleration of gravity is 9.8 m/s 2 . (Let down be negative.) Answer in units of m/s.


1
Expert's answer
2021-09-21T07:21:43-0400

We can find the velocity of the tennis ball when it hits the ground from the kinematic equation:


"v^2=v_0^2+2gs,""v=\sqrt{2gs}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot1.18\ m}=4.81\ \dfrac{m}{s}."